OpenStudy (anonymous):

6. –8x² – 8x – 2 = 0 (1 point) A.one solution B.two solutions C.no solutions D.infinitely many solutions

OpenStudy (anonymous):

b!!

OpenStudy (anonymous):

there's one solution with a multiplicity of 2

OpenStudy (anonymous):

yea sorry its a!!

OpenStudy (anonymous):

u can use the quadratic formula to solve it!

OpenStudy (anonymous):

can one of you guys tell me why :)

OpenStudy (anonymous):

-2(4x^2+4x+1) a=4; b=4; c=1

OpenStudy (anonymous):

\[-4\pm \sqrt{16-16}/8\]

OpenStudy (anonymous):

-4/8= one solution!!

OpenStudy (anonymous):

yes, the discriminant, b^2-4ac equals 0, so there is one solution with a multiplicity of 2

OpenStudy (anonymous):

either 1 or 2

OpenStudy (anonymous):

hmm

OpenStudy (katrinakaif):

2 solutions since the degree of the polynomial is '2'.

OpenStudy (anonymous):

thanks guy! :)

OpenStudy (anonymous):

As two recommended geniuses are saying 2 i will suggest you listen to them

OpenStudy (anonymous):

A question on that, -8x^2-8x=2 How many solutions?

OpenStudy (anonymous):

Im not sure but a lot of people are saying one.

OpenStudy (anonymous):

There is ONE DISTINCT SOLUTION, but yes, it's a degree two polynomial with 2 solutions (you can also say 1 solution with a multiplicity of 2)

OpenStudy (anonymous):

sure thanks Dockworker

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

np

OpenStudy (anonymous):

can someone help me with another question...

OpenStudy (katrinakaif):

Post it on another thread. Start with "Ask a Question"